OpenStudy (anonymous):

is the answer 1285 or 1258?

OpenStudy (anonymous):

V = 1/2 * 4/3 * pi * r^3 V = 2/3 * 3.14* 8.5^3

OpenStudy (anonymous):

4/3πr^3 4/3π 8.5^3 4/3π 614.13 4.1*614.13 2517.93 cm^3 ------------ 2

OpenStudy (anonymous):

which equation is right?

OpenStudy (anonymous):

1286.22 is the answer for r = 8.5

OpenStudy (anonymous):

u still have some problem ? :)

OpenStudy (anonymous):

I got it under control.